# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v2

@@ 6 Prospects
There are still several pivotal suggestions that merit deeper investigation and exploration by the research community going forward.
@@
(2) Development of Multi-Modal Machine Learning Approaches. Furthermore, the investigation of a broader range of MMML techniques is highly encouraged. On the one hand, for time series data of a single modality, it may be feasible to extract features representative of different modalities from the time series data itself, such as the time domain (origin signal), frequency domain (FFT), PSD, STFT, etc. This approach could lead to the preliminary training of unimodal ML models on each single modality, followed by the exploration of MMML strategies. On the other hand, MMML methodologies necessitate more effective representation learning and information alignment techniques—the former concerning the efficient encoding of single modality data, and the latter focusing on the enhanced analysis and fusion of multimodal information for effective PHM prediction or classification tasks. Although simple concatenation is a common method for data fusion in MMML, emerging fusion techniques, such as attention-based or transformer-based mechanisms [111], deserve further exploration. These advanced methods have the potential to effectively capture implicit feature alignments across modalities and facilitate cross-modal synthesis [71, 119]. Yet, research on MMML within the PHM field remains underexplored, highlighting a significant opportunity for the adaptation of existing methods and the innovation of new approaches to address the unique challenges of PHM.
@@
(3) Further Exploration in Machine Learning Model Interpretability. The adoption of DL in PHM has underscored the need for models that are not only high-performing but also interpretable. Techniques such as advanced data visualization and XAI methods are emerging as key tools in explaining the outputs of ML models [64, 6]. These methods are anticipated to provide industries and academia with clearer insights into the decision-making processes of PHM models, thereby building trust and facilitating more informed decision-making [98]. However, current XAI methods predominantly address the interpretability of models using tabular data, text, and images as inputs, leaving a gap in methods tailored for time series data. Moreover, most of the current interpretability methods are applied to unimodal ML models, and the interpretability of deep MMML models has not been explored. Addressing these gaps presents valuable opportunities for future research, aiming to balance the performance and interpretability of ML models. Such advancements would make ML models more trustworthy and interactive, enhancing their utility for scientists and engineers.
@@
(4) Novel Transfer Learning and Domain Adaptation Techniques Development for Model Robustness, and Generalization. Alongside interpretability, the robustness and generalization of ML models in diverse and unpredictable industrial settings are paramount. Novel approaches in transfer learning and domain adaptation can be further developed to ensure models are resilient to data variability and operational uncertainties and capable of adapting to new, unseen scenarios [8]. This ensures their applicability in a wide range of PHM applications, driving the field towards more efficient, reliable, and transparent systems. Currently, research on transfer learning in the PHM domain predominantly addresses fault diagnosis, with only a few studies exploring prognosis. Looking forward, how to utilize the power of transfer learning for prediction problems is still a critical issue. Additionally, cross-modal transfer learning (CMTL) emerges as a critical area of interest in PHM, aiming to improve the knowledge transfer between distinct domains. Moreover, the challenge of collecting a sufficiently large, labeled dataset is a significant barrier in practical applications. The development of unsupervised and semi-supervised transfer learning techniques may help to address this issue.
(5) Construction of Large Knowledge Library. There have seen increasing applications of ML approaches in PHM, but the knowledge and solutions remain fragmented. Therefore, there is a need to establish a comprehensive large knowledge library that can include open-source industrial sensor data, problem-solving approaches, open-source code, model weights, technical papers, review documents, etc. This dynamic knowledge library could evolve over time, becoming more comprehensive and thereby benefiting the broader PHM community. By consolidating various industrial application information into this knowledge library, researchers can quickly access and utilize existing techniques when encountering familiar problems, thus accelerating the pace of research and enhancing problem-solving efficiency. For instance, Lee’s team proposed to construct a industrial large knowledge library including human-interpretable data and structured machine generated data [59]. The proposed industrial LKL could serve as a domain-specific foundation for training Industrial Large Knowledge Models to solve domain-specific problems.
@@
(6) Potential Utilization of Large Language Models (LLMs) and Industrial Large Knowledge Models (ILKMs). Recent advancements in large language model technologies have shown remarkable abilities in natural language processing and related tasks, hinting at the potential for general artificial intelligence applications [125]. In the PHM domain, leveraging these cutting-edge technologies could lead to groundbreaking changes. Yang et al. introduced a novel benchmark dataset focused on Question Answering (QA) in the industrial domain and proposed a new model interaction paradigm, aimed at enhancing the performance of LLMs in domain-specific QA tasks [116]. This approach signifies a substantial stride in customizing LLMs for more specialized, industry-oriented applications. Meanwhile, Li’s team systematically reviewed the current progress and key components of ChatGPT-like large-scale foundation (LSF) models, and provided a comprehensive guide on adapting these models to meet the specific needs of PHM, underscoring the challenges and opportunities for future development [63]. Moreover, as shown in Figure 6, Lee’s team proposed an Industrial Large Knowledge Model (ILKM) framework that aims to solve complex challenges in intelligent manufacturing by combining LLMs and domain-specific knowledge [59]. In summary, by integrating specialized domain knowledge with LLM technology, ML models could offer enhanced performance in addressing industrial data challenges specific to PHM.
Figure 6: Industrial Large Knowledge Model Framework [59]
@@ 7 Conclusion
In summary, ML gradually becomes a cornerstone in PHM, reflecting the big potential for innovative advancements in future PHM development. This paper serves as a valuable resource for both academic and industry professionals in the PHM domain, offering a unified ML framework in PHM and a comprehensive overview of the current state-of-the-art ML approaches for diagnostics and prognostics of industrial systems using industrial open-source data. Based on two primary research task categories: ”Detection & Diagnosis” and ”Assessment & Prognosis”, we provide a detailed explanation of the problems, tasks, challenges, and relevant ML approaches to each competition. Furthermore, we summarize common challenges, including data-related issues and model-related issues, and analyze the solutions to address these challenges. Moreover, we evaluate the limitations of these PHM data challenge competitions and suggest future directions for the PHM data challenge competition evolution. Finally, we prospect six potential directions in the application of data-driven ML within PHM, encompassing a need for open-source multi-modal datasets, development of MMML approaches, further exploration of ML model interpretability, improving the robustness, and generalization of ML models, constructing a large knowledge library, and harnessing the potential of LLMs and ILKMs.
@@
Table 9: Task Description and Website Links for PHM Data Challenge Competitions
PHM Data Challenge Competition Task Description URL
@@
(1) Estimate crack length https://data.phmsociety.org/2019datachallenge/
@@
(1) Predict RUL of the filtration system http://phmeurope.org/2020/data-challenge-2020
@@
(1) RUL prediction in a fleet of aircraft engines https://data.phmsociety.org/2021-phm-conference-data-challenge/
@@
(1) Multiclass fault classification for rock drills https://data.phmsociety.org/2022-phm-conference-data-challenge/

### v3

@@ 5 Prospects
There are still some research directions that deserve deeper investigation and exploration by the research community going forward.
@@
(2) Development of Multi-Modal Machine Learning Approaches. Furthermore, the investigation of a broader range of MMML techniques is highly encouraged. On the one hand, for time series data of a single modality, it may be feasible to extract features representative of different modalities from the time series data itself, such as the time domain (origin signal), frequency domain (FFT), PSD, STFT, etc. This approach could lead to the preliminary training of unimodal ML models on each single modality, followed by the exploration of MMML strategies. On the other hand, MMML methodologies necessitate more effective representation learning and information alignment techniques—the former concerning the efficient encoding of single modality data, and the latter focusing on the enhanced analysis and fusion of multimodal information for effective PHM prediction or classification tasks. Although simple concatenation is a common method for data fusion in MMML, emerging fusion techniques, such as attention based or transformer based mechanisms [117], deserve further exploration. These advanced methods have the potential to effectively capture implicit feature alignments across modalities and facilitate cross-modal synthesis [76, 126]. Yet, research on MMML within the PHM field remains underexplored, highlighting a significant opportunity to explore.
@@
(3) Further Exploration in Machine Learning Model Interpretability. The adoption of DL in PHM has underscored the need for models that are not only high-performing but also interpretable. Techniques such as advanced data visualization and XAI methods are emerging as key tools in explaining the outputs of ML models [68, 6]. These methods are anticipated to provide industries and academia with clearer insights into the decision-making processes of PHM models, thereby building trust and facilitating more informed decision-making [104]. However, current XAI methods predominantly address the interpretability of models using tabular data, text, and images as inputs, leaving a gap in methods tailored for time series data. Moreover, most of the current interpretability methods are applied to unimodal ML models, and the interpretability of MMML models has not been explored. Addressing these gaps can help to balance the performance and interpretability of ML models.
@@
(4) Novel Transfer Learning and Domain Adaptation Techniques Development for Model Robustness, and Generalization. Alongside interpretability, the robustness and generalization of ML models are also important. Novel approaches in transfer learning and domain adaptation can be further developed to ensure models are resilient to data variability and operational uncertainties and capable of adapting to new, unseen scenarios [8]. Currently, research on transfer learning in the PHM domain predominantly addresses fault diagnosis, with only a few studies exploring prognosis. Looking forward, how to utilize the power of transfer learning for prediction problems is still a critical issue. Additionally, cross-modal transfer learning (CMTL) emerges as a critical area of interest in PHM, aiming to improve the knowledge transfer between distinct domains. Moreover, the challenge of collecting a sufficiently large, labeled dataset is a significant barrier in practical applications. The development of unsupervised and semi-supervised transfer learning techniques may help to address this issue.
@@
(5) Potential Utilization of Large Language Models (LLMs) and Industrial Large Knowledge Models (ILKMs). Recent advancements in large language model technologies have shown remarkable abilities in natural language processing and related tasks, hinting at the potential for general artificial intelligence applications [130]. Leveraging these cutting-edge technologies could lead to new changes in PHM domain. Yang et al. introduced a novel benchmark dataset focused on Question Answering (QA) in the industrial domain and proposed a new model interaction paradigm, aimed at enhancing the performance of LLMs in domain-specific QA tasks [123]. This approach signifies a substantial stride in customizing LLMs for more specialized, industry-oriented applications. Meanwhile, Li’s team systematically reviewed the current progress and key components of ChatGPT-like large-scale foundation (LSF) models, and provided a comprehensive guide on adapting these models to meet the specific needs of PHM, underscoring the challenges and opportunities for future development [67]. Moreover, as shown in Figure 6, Lee’s team proposed an Industrial Large Knowledge Model (ILKM) framework that aims to solve complex challenges in intelligent manufacturing by combining LLMs and domain-specific knowledge [63]. Therefore, integrating specialized domain knowledge with LLM technology presents a good opportunity to develop more effective ML models, potentially leading to better solutions for challenges in PHM.
@@ 6 Conclusion
In summary, ML gradually becomes a cornerstone in PHM, reflecting the potential for innovative advancements in future PHM development. This paper serves as a valuable resource for both academic and industry professionals in the PHM domain, offering a unified ML framework in PHM and a comprehensive overview of the current state-of-the-art ML approaches for diagnostics and prognostics of industrial systems using industrial open-source data from recent PHM data challenge. Based on two primary research task categories: ”Detection & Diagnosis” and ”Assessment & Prognosis”, we provide a detailed explanation of the problems, tasks, challenges, and relevant ML approaches to each competition. Furthermore, we summarize common challenges, including data-related and model-related issues, and analyze the solutions to address these challenges. Moreover, we evaluate the limitations of these PHM data challenge competitions and suggest future directions that PHM data challenge competition could focus on. Finally, we prospect five potential research directions in the application of data-driven ML within PHM, encompassing a need for open-source multi-modal datasets, development of MMML approaches, further exploration of ML model interpretability, improving the robustness, and generalization of ML models, and utilization the potential of LLMs and ILKMs.
@@
Table 8: Task Description, References, and Website Links for PHM Data Challenge Competitions
PHM Data Challenge Competition Task Description Reference URL
@@
(1) Estimate crack length
[50, 54]
[127, 92]
https://data.phmsociety.org/2019datachallenge/
@@
(1) Predict RUL of the filtration system
[71, 10, 47, 120]
[64, 30, 113, 46]
http://phmeurope.org/2020/data-challenge-2020
@@
(1) RUL prediction in a fleet of aircraft engines
[73, 103, 25, 26, 104]
[19, 18, 20]
https://data.phmsociety.org/2021-phm-conference-data-challenge/
@@
(1) Multiclass fault classification for rock drills
[86, 53, 79]
[69, 109]
https://data.phmsociety.org/2022-phm-conference-data-challenge/